# Strategic Investment to Mitigate Transition Risks

### v1

@@ 2.5 Optimal Transition Investment Ratio with Government Intervention
Based on the conclusions drawn from the previous subsections, it is evident that companies may exhibit a deficiency in transformational investments in certain scenarios. In such instances, it becomes especially relevant for the government to undertake steps of market adjustments at both macro and micro levels, employing tools such as subsidies or tax penalties.
Subsidies can serve as a powerful incentive for companies by lowering the financial burden associated with making a transition to low-carbon technologies. These subsidies can take various forms, such as direct financial assistance, tax credits, or grants for research and development in green technologies. By reducing the initial costs, subsidies make low-carbon investments more attractive to companies, thereby accelerating the adoption of sustainable practices. For example, a company might receive a tax credit for every unit of renewable energy produced, which would directly offset its production costs and enhance profitability.
On the other hand, tax penalties act as a deterrent against environmentally harmful practices. By imposing higher taxes on carbon emissions or non-compliant activities, governments can create the risk of a financial disincentive for companies to continue with high-carbon operations. This approach compels businesses to internalize the environmental costs of their actions, thereby encouraging them to seek out more sustainable alternatives. For instance, a carbon tax that increases with the level of emissions will drive companies to innovate and invest in cleaner technologies to minimize their tax liabilities.
@@
2.5.1 Model Building
In the extended framework below, we categorize government regulatory tools into two groups: one from a macro-economic perspective and the other from a micro-economic perspective. The government acts as an ‘invisible hand’. We assume that from a macro-economic perspective, the government subsidizes carbon prices that are too high in the market and taxes carbon prices that are too low in the market. On a micro-economic level, if a company surpasses a specified threshold for carbon emissions, the government levies taxes. Conversely, if a company’s total emissions fall below the threshold, the government offers rewards and subsidies to companies, fostering increased enthusiasm for low-carbon transformation initiatives. To introduce the extension model of the the previous framework, additional variables and parameters are introduced to be described in Table 6.
@@ 2.5.1 Model Building
Table 6: Variable and Parameter Description for the Extended Framework
Based on these variables and parameters, the expected total profit of the company after government subsidy or tax can be expressed as:
$\pi^{\prime}=\pi$ $+s_{1}\cdot(\text{ Total Unit Amount })\cdot pr_{1}-q_{1}\cdot(\text{ Total Unit Amount })\cdot\left(1-pr_{1}\right)$ (2.8)
@@
2.5.2 Numerical Study
In this subsection, we expand our analysis to explore the impact of government intervention on the optimal transition investment ratio $(\alpha)$ and profitability of the firm, through a revised objective function that incorporates both subsidies and taxes linked to carbon pricing and emissions. The new objective function, $\pi^{\prime}$, integrates additional terms representing financial incentives and penalties for the firm based on carbon-related metrics.
Our revised framework considers the following variables and parameters: subsidy rates ( $s_{1}$ and $s_{2}$ ) for high carbon prices and low carbon emissions, respectively; tax rates ( $q_{1}$ and $q_{2}$ ) for low carbon prices and high carbon emissions, respectively; and the probabilities ( $pr_{1}$ and $pr_{2}$ ) of these rates being applied based on carbon price and emission thresholds. These factors are critical for determining the cost-benefit landscape of environmental compliance and sustainable investment, as they directly influence the incentives and penalties faced by firms.
The baseline variables from our previous studies are maintained in this study for consistency, but now we add more variables and parameters to the framework. The baseline values for these new variables are shown in Table 7. Then we systematically vary each of the new variables and parameters individually to observe their effects on the optimal value of $\alpha$ and overall profitability for the firm. Graphical representations will be used to illustrate the relationship between these variables and the resulting economic outcomes, providing clear insights into the dynamics at play for the firm under different policy scenarios.
@@ 2.5.2 Numerical Study
From Figure 5 to Figure 10, the left plot illustrates how changes in the key variable affect the optimal transition investment ratio, aiming to maximize profit under each scenario. The right plot shows the corresponding maximum profit achieved under the optimal transition investment as the key variable changes. We will explain each figure in detail below.
Impact of Probability of High Carbon Price Subsidy ($pr_{1}$) on Optimal Alpha and Profit (Figure 5)
As the probability of receiving a subsidy at higher carbon prices ($pr_{1}$) increases, the optimal investment ratio ($\alpha$) for the firm decreases. This suggests that businesses, when expecting financial support through subsidies, feel less pressure to invest heavily in low-carbon technologies. The security provided by expected subsidies reduces the need for extensive investment by the firm in hedging against future regulatory changes or market conditions that favor lower emissions. This also highlights a strategic point for businesses: relying too heavily on subsidies might provide a short-term relief but could risk an insufficient preparation for a future with stricter environmental regulations.
Profit increases alongside higher subsidy probabilities, reinforcing the positive impact of government incentives on financial performance of the firm. For investors, this underscores the appeal of firms with strong expectations of a subsidy support, as subsidies directly enhance profitability of the firm. However, it’s essential to remain cautious—companies that over-rely on subsidies may be less motivated to invest in long-term carbon reduction strategies, posing potential risks to the firm as market dynamics or regulations shift.
@@
As $pr_{2}$ rises, it reflects a higher tolerance from the government for carbon emissions, effectively lowering the threshold required for firms to qualify for subsidies. This shift reduces the incentive for companies to invest in low-carbon transitions, as they can now receive financial support with less stringent emissions standards. Consequently, the optimal value of $\alpha$, which represents the proportion of resources allocated to low-carbon investments, declines. This trend shows that firms are less motivated to prioritize sustainable transitions when subsidies become more accessible without rigorous emission targets.
On the profitability side, a higher $pr_{2}$ initially appears to be advantageous. With reduced pressure to invest in transition efforts, firms can allocate fewer resources to low-carbon initiatives, which boosts short-term profits. However, this profit increase may be deceptive, as it comes at the expense of long-term resilience. Firms that minimize transition investments now could face significant financial risks if stricter carbon policies are introduced in the future. A sudden regulatory shift would leave them underprepared to adapt, potentially leading to substantial losses. This underscores the trade-off between short-term profit gains and the uncertain long-term stability that comes with underinvesting in sustainable practices.
@@
Impact of Tax Rates for Low and High Carbon Emissions ($q_{1}$ and $q_{2}$) on Optimal Alpha and Profit (Figure 7 & Figure 8)
Higher tax rates on low carbon prices ($q_{1}$) or high carbon emissions ($q_{2}$) push the optimal value of $\alpha$ higher, indicating that businesses are incentivized to adopt cleaner technologies more aggressively when facing heavier tax burdens. This reflects the practical use of taxes as a tool for driving environmental sustainability—businesses react to them by increasing investments in green technologies to mitigate the financial impact of taxes. However, businesses must be strategic at the same time: while aggressive adoption of low-carbon technologies can reduce future tax liabilities for the firm, over-investing without aligning with long-term sustainability goals could lead to suboptimal resource allocation.
@@
Impact of Subsidy Rates for High and Low Carbon Scenarios ($s_{1}$ and $s_{2}$) on Optimal Alpha and Profit (Figure 9 & Figure 10)
An increase in subsidy rates for high carbon price ($s_{1}$) or low carbon emission ($s_{2}$) scenarios results in a lower optimal value of $\alpha$, showing that businesses feel less urgency to invest heavily in green technologies when financial incentives make it easier to comply with emission standards. While this may seem beneficial for companies in the short term, there is a risk that such subsidies could slow down the transition to greener technologies if businesses rely too heavily on them without making the necessary levels of investment in long-term sustainability.
@@
Policy Recommendations for Effective Low-Carbon Transitions
Encouraging effective low-carbon transitions requires a sensitive balance between assertive tax policies and a more targeted, limited subsidy approach. The above numerical analysis suggests several important and practical policy recommendations. While subsidies boost short-term profitability, they do not appear to incentivize long-term investment in low-carbon technologies adequately. This is evident from the decreasing trend in the optimal investment ratio ($\alpha$) as subsidy probabilities increase. On the other hand, higher emissions taxes consistently drive higher values of $\alpha$, indicating that tax policies play a more effective role in promoting robust investments in sustainable technologies.
- Expanding on Tax Strategies: A more structured and progressive carbon tax regime could enhance low-carbon transitions more effectively. By gradually increasing the cost of higher emissions, such a policy would not only penalize environmentally harmful practices but also provide businesses with the level of predictability they need in order to plan long-term investments in sustainability. This tax strategy could be further refined by introducing tax incentives for specific technologies that deliver substantial carbon reductions. For businesses, the clear financial penalties for high emissions and rewards for investing in sustainable solutions would create a strong motivation to prioritize clean technology adoption, ensuring a smoother and more predictable transition.
- Reassessing Subsidy Schemes: A more passive and targeted approach to subsidies is recommended. Direct financial incentives should be limited to sectors of the economy where low-carbon alternatives are not yet economically viable or where early-stage technologies need support to scale. This approach minimizes the risk of companies relying on subsidies for profitability without making a genuine progress toward sustainability. It also encourages firms to seek out real, long-term solutions rather than temporary financial relief. Companies should anticipate stricter regulatory standards in the future and take a proactive approach to sustainability. Rather than depending on subsidies alone, businesses are advised to prioritize investments in scalable, low-carbon technologies that align with long-term industry trends. Building internal capabilities for sustainable practices and integrating environmental risk assessments into financial planning can help ensure resilience against future regulatory shifts. Additionally, companies should explore partnerships with other industry players or research institutions to advance low-carbon innovations, as these collaborations can reduce costs and accelerate progress toward sustainability goals.
@@ 3 Conclusions
The study suggests a balanced policy approach—where taxes are used assertively to stimulate investment, and subsidies are applied selectively to avoid dependency. This would encourage corporations to take proactive steps toward sustainability, aligning with long-term environmental goals rather than relying solely on government incentives. For stakeholders, particularly investors, the research underscores the importance of identifying companies that are actively adapting to evolving fiscal and regulatory landscapes. These companies are better positioned to meet future regulatory demands and capitalize on the growing market emphasis on sustainability, offering better long-term returns.
Extending the analytical framework to a multi-period setting under different decarbonization scenarios further reveals how the optimal investment ratio ($\alpha$) evolves over time. While aggressive early decarbonization yields the highest short-term profits for the firm, slower approaches offer better long-term profitability due to the lower upfront costs. The analysis also shows that sustainable investment in low-carbon technologies is beneficial in the long run, particularly in scenarios involving gradual transitions. In conclusion, this study provides practical insights for policymakers and business leaders on how to promote investments in low-carbon technologies. By implementing a balanced mix of taxes and targeted subsidies, governments can effectively encourage corporate sustainability efforts.
While the numerical examples presented in this paper provide valuable insights into the dynamics of decarbonization strategies, there is a noticeably potential to extend the model into a more general framework that relies on abstract variables rather than specific parameter values. One direction for this generalization is to define the production function, cost structure, and investment dynamics in terms of general parameters (e.g., $f(\alpha,k)$ for production efficiency, $g(A)$ for asset allocation) without relying on the sigmoid function or fixed coefficients such as $\beta$ and $k$. By introducing generalized mathematical forms, the model could capture a wider variety of industry-specific or region-specific decarbonization scenarios, making it applicable to diverse contexts.
Another avenue for extension lies in relaxing some of the simplifying assumptions made in this paper. For example:
  1. 1.
    The framework could incorporate multi-factor production functions, such as CES, that allow for varying degrees of substitutability between capital and efficiency. This would provide more flexibility in modeling how firms adapt to changing technological and regulatory environments.
  2. 2.
    Introducing uncertainty variables to account for fluctuations in energy prices, regulatory costs, or technological advancements could make the framework more robust to real-world complexities.
  3. 3.
    Expanding the time horizon to include dynamic decision-making models would allow firms to adjust their investment strategies iteratively in response to external changes.
  4. 4.
    Making a transition from a partial equilibrium framework to a general equilibrium framework to capture first-order welfare effects that are important for drawing policy recommendations.
Finally, transitioning to a more generalized theoretical framework would facilitate the development of closed-form solutions or approximate analytical results for key metrics, such as optimal investment ratios or profit trajectories, under a range of assumptions. This would reduce reliance on numerical simulations while retaining the core insights of the model. These extensions would not only enhance the versatility of the framework but also deepen its relevance for policymakers, firms, and researchers exploring sustainable transitions in diverse industries.

### v2

@@ 2.5 Optimal Transition Investment Ratio with Government Intervention
From the conclusions drawn from the previous subsections, it is evident that companies may exhibit a deficiency in transformational investments in certain scenarios. In such instances, it becomes especially relevant for the government to undertake steps of market adjustments at both macro and micro levels, employing tools such as subsidies or tax penalties.
Subsidies can serve as a useful incentive for companies by lowering the financial burden associated with making a transition to low-carbon technologies. These subsidies can take various forms, such as direct financial assistance, tax credits, or grants for research and development in green technologies. By reducing the initial costs, subsidies make low-carbon investments more attractive to companies, thereby accelerating the adoption of sustainable practices. For example, a company might receive a tax credit for every unit of renewable energy produced, which would directly offset its production costs and enhance profitability.
On the other hand, tax penalties can act as a deterrent against environmentally harmful practices. By imposing higher taxes on carbon emissions or non-compliant activities, governments can create the risk of a financial disincentive for companies to continue with high-carbon operations. This approach compels businesses to internalize the environmental costs of their actions, thereby encouraging them to seek out more sustainable alternatives. For instance, a carbon tax that increases with the level of emissions will drive companies to innovate and invest in cleaner technologies to minimize their tax liabilities.
@@
In the extended model to be presented below, we categorize government regulatory tools into two groups: one from a macro-economic perspective and the other from a micro-economic perspective. The government acts as an ‘invisible hand’. We assume that from a macro-economic perspective, the government subsidizes carbon prices that are too high in the market and taxes carbon prices that are too low in the market. At a micro-economic level, if a company surpasses a specified threshold for carbon emissions, the government levies taxes. Conversely, if a company’s total emissions fall below the threshold, the government offers rewards and subsidies to companies, fostering increased enthusiasm for low-carbon transformation initiatives. To introduce the extension of the the previous model, additional variables are introduced in Table 6.
@@
Table 6: Variable Description for the Extended Framework
Based on these variables, the expected total profit of the company after government subsidy or tax can be expressed as:
[…]
@@
Next, we expand our numerical analysis to explore the impact of government intervention on the optimal transition investment ratio $(\alpha)$ and profitability of the firm, through a revised objective function that incorporates both subsidies and taxes linked to carbon pricing and emissions. The new objective function, $\pi^{\prime}$, integrates additional terms representing financial incentives and penalties for the firm based on carbon-related metrics.
Our expanded model considers the following variables: subsidy rates ( $s_{1}$ and $s_{2}$ ) for high carbon prices and low carbon emissions, respectively; tax rates ( $q_{1}$ and $q_{2}$ ) for low carbon prices and high carbon emissions, respectively; and the probabilities ( $pr_{1}$ and $pr_{2}$ ) of these rates being applied based on carbon price and emission thresholds. These factors are critical for determining the cost-benefit landscape of environmental compliance and sustainable investment, as they directly influence the incentives and penalties faced by firms.
The baseline variables from our previous studies are maintained in this study for consistency, but now we add more variables to the model. The baseline values for these new variables are shown in Table 7. Then we systematically vary each of the new variables individually to observe their effects on the optimal value of $\alpha$ and overall profitability for the firm. Graphical representations will be used to illustrate the relationship between these variables and the resulting economic outcomes, providing clear insights into the dynamics at play for the firm under different policy scenarios.
@@
From Figure 5 to Figure 10, the left plot illustrates how changes in the key variable affect the optimal transition investment ratio with the goal of maximizing profit under each scenario. The right plot shows the corresponding maximum profit achieved by the firm under the optimal transition investment as the key variable changes. We will explain each figure in detail below.
Impact of Probability of High Carbon Price Subsidy ($pr_{1}$) on Optimal Alpha and Profit (Figure 5). As the probability of receiving a subsidy at higher carbon prices ($pr_{1}$) increases, the optimal investment ratio ($\alpha$) for the firm decreases. This suggests that businesses, when expecting financial support through subsidies, feel less pressure to invest heavily in low-carbon technologies. The security provided by expected subsidies reduces the need for extensive investment by the firm in hedging against future regulatory changes or market conditions that favor lower emissions. This also highlights a strategic point for businesses: relying too heavily on subsidies might provide a short-term relief but could risk an insufficient preparation for a future with stricter environmental regulations.
Profit increases alongside higher subsidy probabilities, reinforcing the positive impact of government incentives on financial performance of the firm. For investors, this underscores the appeal of firms with strong expectations of a subsidy support, as subsidies directly enhance profitability of the firm. However, it is essential to remain cautious – companies that over-rely on subsidies may be less motivated to invest in long-term carbon reduction strategies, posing potential risks to the firm as market dynamics or regulations shift.
@@
A rising $pr_{2}$ reflects a higher tolerance from the government for carbon emissions, effectively lowering the threshold required for firms to qualify for subsidies. This shift reduces the incentive for companies to invest in low-carbon transitions, as they can now receive financial support with less stringent emissions standards. Consequently, the optimal value of $\alpha$, which represents the proportion of resources allocated to low-carbon investments, declines. This trend shows that firms are less motivated to prioritize sustainable transitions when subsidies become more accessible without rigorous emission targets.
On the profitability side, a higher $pr_{2}$ initially appears to be advantageous. With a reduced amount of pressure to invest in transition efforts, firms can allocate fewer resources to low-carbon initiatives, which boosts short-term profits. However, this profit increase may be deceptive, as it comes at the expense of long-term resilience. Firms that minimize transition investments now could face significant financial risks if stricter carbon policies are introduced in the future. A sudden regulatory shift would leave them underprepared to adapt, potentially leading to substantial losses. This underscores a trade-off between short-term profit gains and the uncertain long-term stability that comes with underinvesting in sustainable practices.
@@
Impact of Tax Rates for Low and High Carbon Emissions ($q_{1}$ and $q_{2}$) on Optimal Alpha and Profit (Figure 7 & Figure 8). Higher tax rates on low carbon prices ($q_{1}$) or high carbon emissions ($q_{2}$) push the optimal value of $\alpha$ higher, indicating that businesses are incentivized to adopt cleaner technologies more aggressively when facing heavier tax burdens. This reflects the practical use of taxes as a tool for driving environmental sustainability—businesses react to them by increasing investments in green technologies to mitigate the financial impact of taxes. However, businesses must be strategic at the same time: while aggressive adoption of low-carbon technologies can reduce future tax liabilities for the firm, over-investing without aligning with long-term sustainability goals could lead to a suboptimal resource allocation.
@@
Impact of Subsidy Rates for High and Low Carbon Scenarios ($s_{1}$ and $s_{2}$) on Optimal Alpha and Profit (Figure 9 & Figure 10). An increase in subsidy rates for high carbon price ($s_{1}$) or low carbon emission ($s_{2}$) scenarios results in a lower optimal value of $\alpha$, showing that businesses feel less urgency to invest heavily in green technologies when financial incentives make it easier to comply with emission standards. While this may seem beneficial for companies in the short term, there is a risk that such subsidies could slow down the transition to greener technologies if businesses rely too heavily on them without making the necessary levels of investment in long-term sustainability.
@@
Policy Recommendations for Effective Low-Carbon Transitions. Encouraging effective low-carbon transitions requires a sensitive balance between assertive tax policies and a more targeted, limited subsidy approach. The above numerical analysis suggests several important and practical policy recommendations. While subsidies boost short-term profitability, they do not appear to incentivize long-term investment in low-carbon technologies adequately. This is evident from the decreasing trend in the optimal investment ratio ($\alpha$) as subsidy probabilities increase. On the other hand, higher emissions taxes consistently drive higher values of $\alpha$, indicating that tax policies play a more effective role in promoting robust investments in sustainable technologies.
Expanding on Tax Strategies. A more structured and progressive carbon tax regime could enhance low-carbon transitions more effectively. By gradually increasing the cost of higher emissions, such a policy would not only penalize environmentally harmful practices but also provide businesses with the level of predictability they need in order to plan long-term investments in sustainability. This tax strategy could be further refined by introducing tax incentives for specific technologies that deliver substantial carbon reductions. For businesses, the clear financial penalties for high emissions and rewards for investing in sustainable solutions would create a strong motivation to prioritize clean technology adoption, ensuring a smoother and more predictable transition.
Reassessing Subsidy Schemes. A more passive and targeted approach to subsidies is recommended. Direct financial incentives should be limited to sectors of the economy where low-carbon alternatives are not yet economically viable or where early-stage technologies need support to scale. This approach minimizes the risk of companies relying on subsidies for profitability without making a genuine progress toward sustainability. It also encourages firms to seek out real, long-term solutions rather than temporary financial relief. Companies should anticipate stricter regulatory standards in the future and take a proactive approach to sustainability. Rather than depending on subsidies alone, businesses are advised to prioritize investments in scalable, low-carbon technologies that align with long-term industry trends. Building internal capabilities for sustainable practices and integrating environmental risk assessments into financial planning can help ensure resilience against future regulatory shifts. Additionally, companies should explore partnerships with other industry players or research institutions to advance low-carbon innovations, as these collaborations can reduce costs and accelerate progress toward sustainability goals.
@@ 3 Conclusions
The study suggests a balanced policy approach—where taxes can be used assertively to stimulate investment, and subsidies can be applied selectively to avoid dependency. This would encourage corporations to take proactive steps toward sustainability, aligning with long-term environmental goals rather than relying solely on government incentives. For stakeholders, particularly investors, the research underscores the importance of identifying companies that are actively adapting to evolving fiscal and regulatory landscapes. These companies are better positioned to meet future regulatory demands and capitalize on the growing market emphasis on sustainability, offering better long-term returns.
Extending the one-period model to a multi-period setting under different decarbonization scenarios further reveals how the optimal investment ratio ($\alpha$) evolves over time. While aggressive early decarbonization yields the highest short-term profits for the firm, slower approaches offer better long-term profitability due to the lower upfront costs. The analysis also shows that sustainable investment in low-carbon technologies is beneficial in the long run, particularly in scenarios involving gradual transitions. In conclusion, this study provides practical insights for policymakers and business leaders on how to promote investments in low-carbon technologies. By implementing a balanced mix of taxes and targeted subsidies, governments can effectively encourage corporate sustainability efforts.